Does any solid Ag2CrO4 form when 8.0 10-4 g of AgNO3 is dissolved in 15 mL of 6.6 10-5 M K2CrO4? Assume no volume change. (NOTE: Only ONE submission is allowed for the first part of this question.)

Yes, precipitate will form.

No, precipitate will not form.

Qsp?

Ksp?
Answer #1

Formation of teh percipitate of the compound in any reaction is dependent on the comparison of the ionic product Qsp and the solubility product Ksp.

If the value of inoic product Qsp is greater than the solubility product Ksp then there is formation of the pericipite.

If the value of inoic product Qsp is smaller than the solubility product Ksp then there is formation of the pericipitate do not takes place.

THE VALUE OF Ksp, FOR THE Ag2CrO4 = 1.2*10^{-12}

now the calculation part is :

given 8x10 gm of Agnoz is dissolved now molecular weight of AgNO : 169.879/mo So no. of moles of AgNo= - Given mas . molar ma

now calculated value of ionic product Qsp is = 6.5*10^{-12} , which is greater than the Ksp so ppt formation will occurs.